### v1

# Stark localization near Aubry-André criticality

Abstract
In this work we investigate the Stark localization near the Aubry-André (AA) critical point. We study system-dependent parameters, such as localization length, inverse participation ratio (IPR), and energy gap between the ground and first excited state, for characterizing the localization-delocalization transition. We show that the scaling exponents possessed by these key descriptors of localization are quite different from that of a pure AA model or Stark model. Near the critical point of the AA model, inducing Stark field of strength $h$, the localization length $\zeta$ scales as $\zeta\propto h^{-\nu}$ with $\nu\approx 0.29$ which is different than both the pure AA model ($\nu=1$) and Stark model ($\nu\approx 0.33$). The IPR scales as IPR $\propto h^{s}$ with $s\approx 0.096$ which is again significantly different than both the pure AA model ($s\approx 0.33$) and Stark model ($s\approx 0.33$). The energy gap, $\Delta$, scales as $E\propto h^{\nu z}$, where $z\approx 2.37$ which is however same as the pure AA model.
I Introduction
Randomness is an inherent property of nature whose influence on any physical system often defies our intuition and yields non-trivial effects. Consequently, disordered systems have long been a focal point of significant research endeavors Binder86 ; Belitz05 ; Das08 ; Alloul09 ; Konar22 ; Ghosh21 . Within the realm of quantum mechanics, the interplay of disorder with fundamental principles gives rise to a myriad of noteworthy phenomena, including the emergence of novel quantum phases, order-from-disorder phenomena Aizenman89 ; Wehr06 ; Bera14 ; Bera16 ; Bera17 , the enigmatic high-$T_{c}$ superconductivity Auerbach94 , and the phenomenon of localization Anderson58 ; Abrahams79 ; Evers08 ; Nandkishore15 ; Abanin19 ; Aitijhaya23 . Among these phenomena, disorder-induced Anderson localization occupies a prominent position Yamilov23 ; Lagendijk09 . This well-established concept elucidates how the presence of randomness can hinder the diffusion of waves within a system, showcasing its pivotal role in modern condensed matter physics. In Anderson model the disorder is truly random and uncorrelated. Whereas some disorders are not random, rather there is some correlation. Quasi-periodicity is the candidate that qualifies for this correlated disorder.
In the recent past and even nowadays quasiperiodic systems Sinha19 ; Wei19 are widely studied as it captures very exotic features of various novel phases of matter and supports phase transition between those states Chowdhury86 ; Mezard87 ; Sachdev99 ; Yao14 ; Zuniga13 . Among these models, the Aubry-André (AA) Aubry80 ; Sahoo24 model is of central interest of researchers for its distinct characteristics in the realm of phase transitions Zanardi2006 ; Mondal23 . One of the distinguishing features of the AA model is its self-dual symmetry, evident between the Hamiltonian formulations in momentum and position spaces. This symmetry manifests in an energy-independent localization-delocalization transition Michal14 ; Iyer13 occurring at a finite modulation strength. This model has been extensively explored across various contexts, including the elucidation of the Hofstadter butterfly structure of the energy spectrum under controlled parameters Hofstadter76 , investigations into transport phenomena Purkayastha18 ; Sutradhar19 , analysis of mobility edges Saha16 ; Ganeshan15 ; Ganeshan1502 , critical behavior, and exploration of topological phases DeGottardi13 ; Cai13 ; Fraxanet3 , many-body localization Gu14 ; Zhang08 ; Zhang09 , dynamical phase transition Modak21 , in context of open quantum system Carmichael93 ; Xu16 ; Kawabata17 ; Lee19 , and for its applications in quantum technology Sahoo24 . The AA model has been realized using ultracold atoms in an incommensurate optical lattice Makris08 ; Klaiman08 ; Ruter10 and photonic lattices Roati08 .
The AA model is a tight binding model with nearest neighbour tunnelling in the presence of sinusoidal quasi-periodic potential. When the strength of the potential becomes twice the tunneling parameter, in the thermodynamic limit, the system experiences a phase transition from a delocalized phase to a localized phase. On the other hand, in a tight binding model, if a gradient field is induced across the lattice making the on-site energies off-resonant, the tunnelling rate gets suppressed resulting in the localization of wavefunction in space in absence of disorder. In the thermodynamic limit, this localization, known as Stark localization Schulz19 , takes place at zero field limit. A substantial amount of research has been conducted for understanding Stark localization phenomena in context of single-particle Wannier60 ; Fukuyama73 ; Holthaus95 ; Kolovsky03 ; Kolovsky08 ; Kolovsky13 ; Nieuwenburg19 and many-body Jiang23 ; Zhang21 ; Yao21 ; Wei22 ; Doggen21 systems, and for finding applications Morong21 ; Preiss15 ; Karamlou22 ; He23 ; Kohlert21 by exploiting it.
All these are quantum phase transitions that manifest at zero temperature when the control parameter of a quantum system’s Hamiltonian is finely tuned to a critical value, referred to as the quantum critical point (QCP). At this juncture, the system exhibits scaling and universality Hikami81 ; Altland97 ; Wang21 ; Luo21 ; LuoR22 ; Luo22 ; Bu22 , meaning that the equilibrium properties of certain physical observables near the QCP can be succinctly described by a handful of critical exponents. Extracting these critical exponents necessitates an understanding of various physical quantities  Mishra16 ; Sadhukhan15 ; Sadhukhan16 , such as the localization length, Inverse Participation Ratio (IPR), energy gap, and fidelity susceptibility etc. These quantities provide essential insights into the behavior of the system near the QCP and facilitate the characterization of its critical behavior. In this work, we consider a single particle tight-binding model with nearest-neighbour hopping. The system is prepared at the QCP of the AA model. Then a Stark potential is introduced for studying the scaling properties of localization length, IPR, and energy gap between the ground and the first excited state.
Following an introductory overview in Section I, Section II presents the model of our system. In Section III, defining the characteristic quantities relevant to the localization-delocalization transition, the behavior of the corresponding scaling exponents at the criticality of the AA model is studied via the cost function approach. In Section IV the same scaling exponents are computed in the vicinity of the critical point of the AA model. In Section V, we explore the evolution of the scaling exponent of IPR as we move away from the critical region of the AA model and additionally, we discuss how the corresponding structures of the wave function change. Then in Section VI, we introduce a hybrid scaling exponent that is related to the other scaling exponents of the pure AA and Stark model. Finally, a conclusion is given in Section VII.
II Model
The Hamiltonian of the Aubry-André model with a linear gradient field across the lattice, is
$\hat{H}=$ $-\;J$ $\sum_{i}^{L}(\hat{c}^{\dagger}_{i}\hat{c}_{i+1}+\mathrm{h.c.})+h\sum_{i}^{L}i\hat{c}^{\dagger}_{i}\hat{c}_{i}$ (1)
$+$ $(2J+\delta)\sum_{i}^{L}\cos[2\pi(i\omega+\phi)]\hat{c}^{\dagger}_{i}\hat{c}_{i}$
where, $\hat{c}^{\dagger}_{i}(\hat{c}_{i})$ is the fermionic creation (annihilation) operator at the $i^{\mathrm{th}}$ site, $(2J+\delta)$ determines the amplitude of the quasiperiodic modulation, $\omega$ is an irrational number has a special form of $\omega=F_{n}/F_{n+1}$ where $F_{n}$ are the Fibonacci numbers and system size $L$ has been chosen as $L=F_{n+1}$, $\phi$ is a random number resulting in some random phase of quasiperiodic potential, drawn from a uniform distribution in $[0,1)$, and $h$ quantifies the magnitude of the Stark field. The boundary condition used in this work is the open boundary condition (OBC). In the thermodynamic limit (i.e., $L\to\infty$) for the absence of the second term, i.e., $h=0$, one finds that all the eigenstates are localized for $\delta>0$ and all the eigenstates are delocalized for $\delta<0$. On the other hand, without the last term i.e., $\delta=-2J$, in the thermodynamic limit, the system is localized for $h\to 0$.
III Scaling Analysis at the AA Criticality
We start our study of the behaviour of the scaling exponents right at the AA criticality, i.e., $\delta=0$. In a finite system, at this point, the wavefunction of the particle is neither localized nor delocalized, rather there happens to be a continuous phase transition. To quantify the localized or delocalized nature of the wavefunction, we use the observable localization length, $\zeta$, defined as,
$\zeta=\sqrt{\sum_{i}^{L}(i-i_{c})^{2}\;p_{i}},$ (2)
where $i$ denotes the lattice site, $p_{i}$ is the single-particle probability density at site $i$ and $i_{c}$ is the localization center having the expression $i_{c}=\sum_{i=1}^{L}ip_{i}$. Expanding the $n^{\mathrm{th}}$ normalised eigenstate of the system, $|\psi^{n}\rangle$, in terms of the single-particle computational basis, $|i\rangle$, such that $|\psi_{n}\rangle=\sum_{i}c^{(i)}_{n}|i\rangle$, $p_{i}$ is given by $p_{i}=|\langle i|\psi^{n}\rangle|^{2}=|c^{(i)}_{n}|^{2}$.
Figure 1: Scaling Analysis at $\delta=0$: (a1) presents localization length $\zeta$ of the ground state against Stark field strength $h$ for different system sizes, $L=144$ (blue circle), $L=233$ (violet square), $L=377$ (indigo diamond), $L=610$ (maroon triangle up) and $L=987$ (orange triangle left). (a2) shows a collapse plot for the localization length with the value of scaling exponent $\nu=0.29$. (b1) displays the IPR versus $h$ for various values of $L$ which is aforementioned in (a1). (b2) demonstrates the collapse plot of IPR using the value of scaling exponent $s=0.096$ and $\nu=0.29$. (c1) shows the energy gap $\Delta E$ between the ground state and the first excited state as a function of $h$ for different $L$ values as mentioned previously. (c2) demonstrates the collapse plot of $\Delta E$ using the scaling exponent $z=2.37$. we have taken $\phi$ average for 5000 random choices in the range $[0,1)$.
If $g$ is the control parameter in a certain localization-delocalization transition and $g_{c}$ is the critical point then in the thermodynamic limit, near this criticality, the localization length scales as,
$\zeta\propto|g-g_{c}|^{-\nu},$ (3)
where $\nu$ is the scaling exponent. For the pure AA model, $|g-g_{c}|=\delta$ and $\nu=1$; for the pure Stark model $|g-g_{c}|=h$ and $\nu\approx 0.33$. Near criticality, the localization length scales with system size as,
$\mathrm{\zeta}\propto L^{-1}.$ (4)
Figure 2: Scaling Analysis for $\delta L^{1/\nu}=1$: (a1) presents localization length $\zeta$ against the Stark field of strength $h$ for different system sizes, $L=144$ (blue circle), $L=233$ (violet square), $L=377$ (indigo diamond), $L=610$ (maroon triangle up), $L=987$ (orange triangle left). (a2) shows the corresponding collapse plot for a value of the scaling exponent $\nu=0.29$. (a3) displays the IPR versus $h$ for various values of $L$ which is aforementioned in (a1). (a4) demonstrates the collapse plot of IPR using the scaling exponent value $s=0.096$. Scaling Analysis for $\delta L^{1/\nu}=-1$: (b1) presents localization length $\zeta$ against stark potential strength $h$ for different system sizes, $L$ with the same line-point scheme as before. (b2) shows the collapse plot for a value of scaling exponent $\nu=0.29$. (b3) displays the IPR versus $h$ for various values of $L$ which is aforementioned in (a1). (b4) demonstrates the collapse plot of IPR using the scaling exponent value $s=0.096$.
We study the dependence of the localization length $\zeta$, of the ground state, on the control parameter $h$ for different system sizes $L$. For localization length, we adopt the following scaling ansatz,
$\zeta/L=f_{1}(hL^{1/\nu})$ (5)
where $f_{1}[\cdot]$ is an arbitrary function. This scaling ansatz is confirmed through the methodology of data collapse. The concept behind data collapse is as follows: We collect the data $\zeta/L$ versus $hL^{1/\nu}$ for various system sizes $L$, with a specific value of $\nu$, for which all the datasets collapse perfectly. Initially, by some trial and error, we manually find out a rough range of the scaling exponent for which the data collapse is quite well. Furthermore, for a better estimation of the scaling exponent we employ the cost function approach Suntajs20 , where the cost function $C_{Q}$ is defined as,
$C_{Q}=\frac{\sum_{i=1}^{N_{p}-1}|Q_{i+1}-Q_{i}|}{\mathrm{max}\{Q_{i}\}-\mathrm{min}\{Q_{i}\}}-1,$ (6)
where $\{Q_{i}\}$ is the dataset for different values of $h$ and $L$ and $N_{p}$ is the total number of data points in the dataset. Sorting all $N_{p}$ values of $|Q_{i}|$ according to the increasing value of $L\,\mathrm{sgn}[h-h_{c}]\,[h-h_{c}]^{\nu}$, the best scaling exponent can be found for the minimum value of cost function $C_{Q}$. Ideally, the perfect data collapse is found for $C_{Q}=0$, i.e., $\sum_{i=1}^{N_{p}-1}|Q_{i+1}-Q_{i}|=\mathrm{max}\{Q_{i}\}-\mathrm{min}\{Q_{i}\}.$
In Fig. 1(a1), we depicted $\zeta$ against $h$ for various values of $L$. This figure illustrates that for finite-sized systems, an initial flat region exists wherein the corresponding wave function exhibits an extended nature, indicating that the ground state of the Hamiltonian for $\delta=0$ is equally probable across this phase. Beyond a certain threshold of $h$, the localization length exhibits weak dependence on the system size, leading to the localization of the wavefunction. To derive the scaling exponent $\nu$ from Eq. (3) with $|g-g_{c}|=h$, we demonstrate a collapse plot in Fig. 1(a2) utilizing the function specified previously in Eq. (5). The best collapse is obtained for $\nu=0.29(1)$. The cost function approach further confirms this optimal value of the scaling exponent $\nu$.
The next observable that can capture the delocalization-localization transition is the IPR, defined as,
$\mathrm{IPR}=\sum_{i=1}^{L}p_{i}^{2}.$ (7)
Near criticality, in the thermodynamic limit, the IPR scales as,
$\mathrm{IPR}\propto|g-g_{c}|^{s},$ (8)
where $s$ is another scaling exponent. For the pure AA model, $|g-g_{c}|=\delta$ and $s\approx 0.33$; for the pure Stark model $|g-g_{c}|=h$ and $s\approx 0.33$. On the other hand, near criticality, IPR scales with system size $L$ as,
$\mathrm{IPR}\propto L^{-s/\nu}.$ (9)
So for IPR, the scaling ansatz we use is,
$\mathrm{IPR}=L^{-s/\nu}f_{2}(hL^{1/\nu}),$ (10)
where $f_{2}[\cdot]$ is another arbitrary function.
The variation in the value of IPR with the increase in $h$ for different system sizes $L$ is depicted in Fig. 1(b1). In this figure, upto a certain region of $h$, the IPR values appear flat, indicating the presence of a delocalized phase. Beyond this region, the IPR becomes independent of the system size, characterizing the localized phase. We employ the cost function approach which suggests $s=0.096(1)$. Fig. 1(b2) presents the best collapse plot of the IPR using $s=0.096$. Noticeably, in this case, $s/\nu\approx 0.33$ which is the same as the pure AA model demonstrating the fact that at the criticality the scaling of IPR with the system size is the same for both of these two models.
Figure 3: Scaling Analysis for $\delta=-0.1$: (a1) presents localization length $\zeta$ against stark potential strength $h$ for different system size $L$, with the same line-point scheme. (a2) shows the collapse plot for a value of the scaling exponent $\nu=0.33$. (b1) displays the IPR versus $h$ for various values of $L$. (b2) demonstrates the collapse plot of IPR with $s=0.33$. (c1) depicts the variation of $\Delta E$ as a function of $h$ for different $L$’s. (c2) demonstrates the collapse plot of $\Delta E$ using the scaling exponent $z=2$ as before.
Finally, to elucidate the localization transition, we examine the energy gap between the ground state and the first excited energy, denoted as $\Delta E$. In the thermodynamic limit, the energy gap scales with the control parameter as,
$\Delta E\propto|g-g_{c}|^{\nu z}.$ (11)
For the pure AA model $z=2.37$ and for the pure Stark model $z=2$. To perform the finite size scaling for $\Delta E$, we use the scaling ansatz:
$\Delta E=L^{-z}f_{3}(hL^{1/\nu}),$ (12)
where $f_{3}[\cdot]$ is another arbitrary function.
Fig. 1(c1) shows, the energy gap $\Delta E$ against the Stark field strength $h$ for different system sizes. In this case, also there is an initial flat region showing a very narrow energy gap in the delocalized phase. Beyond a certain strength of Stark field $h$, $\Delta E$ becomes independent of the system size and the energy gap gets increased with an increase in the field strength. The scaling exponent $z$ is determined through data collapse yielding a critical exponent of $z=2.37(1)$. It is also the same as the pure AA model. The collapse plot corresponding to $\Delta E$ is presented in Fig. 1(c2).
Figure 4: (a1) presents $s/\nu$ for different values of $\delta$. For $\delta=0$ the $s/\nu$ value is $0.33$ which is the same as that of the pure AA model and this value continuously increases to $1$, which is the same as the pure Stark model, as one moves away from the AA criticality with negative $\delta$ values. (a2) presents the plot of $z$ for different values of $\delta$. In this case also, there is a continuous change in the value of $s/\nu$ from 2.37, which is the same as the pure AA model, to 2, which is for the pure Stark model, as one moves away from the criticality.
IV Scaling Properties near AA criticality
In the previous section, we have done the scaling analysis at the AA criticality, i.e., $\delta=0$. Now we turn our attention to the near criticality region of the AA model. Now as $\delta\neq 0$ one can expect that the previous scaling ansatz will not work properly and some modifications are needed.
We start with the scaling of localization length. In the presence of both $\delta$ and $h$, we use the following ansatz for the scaling of localization length,
$\zeta/L=f_{4}(\delta L^{1/\nu_{\delta}},hL^{1/\nu}),$ (13)
where $\nu_{\delta}$ is the scaling exponent for localization length in the case of a pure AA model. The novelty of this ansatz is that when $\delta=0$, it coincides with the ansatz given in Eq. (5), and for $h=0$, it becomes the scaling ansatz for pure AA model, i.e., $\mathrm{\zeta}\propto\delta^{-\nu_{\delta}}$, in the thermodynamic limit.
In the same spirit, we define the scaling ansatz for IPR as,
$\mathrm{IPR}=L^{-s_{\delta}/\nu_{\delta}}f_{5}(\delta L^{1/\nu_{\delta}},hL^{1/\nu})$ (14)
where $s_{\delta}$ is the scaling exponent of IPR in the pure AA model ($h=0$), i.e., $\mathrm{IPR}\propto\delta^{s_{\delta}}$, in the thermodynamic limit. When $\delta=0$, this scaling ansatz takes the similar form used in the previous section for the collapse as we have found that the $s/\nu$ value is the same for both the pure AA model and our model under consideration. Similarly, for $\Delta E$ it will take the form,
$\Delta E=L^{-z_{\delta}}f_{6}(\delta L^{1/\nu_{\delta}},hL^{1/\nu}),$ (15)
where $z_{\delta}$ appears in the scaling exponent of $\Delta E$ in the pure AA model, i.e., $\Delta E\propto\delta^{\nu_{\delta}z_{\delta}}$, in the thermodynamic limit, again this works due to the same value of $z$ in both models.
Figure 5: Ground State Wavefunction $\psi_{0}$: (a1) describes the ground state wavefunction for the pure AA model, where $h=0$ and $\delta=0.4J$. (a2) presents the ground state wavefunction for the pure stark model, where $\delta=-2J$ and $h=10^{-4}$. (a3) presents the ground state wavefunction for $\delta=0$ with $h=10^{-4}$ and (a4) presents the wave function for $\delta=-0.1$ and $h=10^{-4}$. All plots are done for system size 610 and for the AA potential, the phase, $\phi=0$.
To examine the validation of the above scaling ansatz, we now turn our attention to delve into scaling properties of $\zeta$, IPR, and $\Delta E$ near criticality for $\delta L^{1/\nu}=1$ and $\delta L^{1/\nu}=-1$. Upon taking $\delta$ in this fashion for different $L$, we expect that our previous critical exponents would work as we are still in the vicinity of the AA critical point. In Fig. 2 both these cases are shown and we have numerically found that the collapse still happens for the same set of scaling exponents as found in the case of $\delta=0$.
V Scaling Analysis at Far from the Aubry-André Criticality
Now we focus on the scaling analysis at comparatively far away from the AA criticality. Far from the AA criticality, the wavefunction is supposed to be delocalized as the effect of AA potential is faint at that region. In order to validate this understanding we perform a scaling analysis for a fixed value of $\delta<0$, specifically $\delta=-0.1$, as illustrated in Fig. 3.
In Fig. 3(a1), the localization length ($\zeta$) is plotted against $h$, while in Fig. 3(a2), we depict $\zeta/L$ against $hL^{1/\nu}$, with a specific value of $\nu=0.33$. Remarkably, all datasets collapse onto a single curve for this value of $\nu$. This observation suggests that in the as we move away from the AA criticality the scaling nature replicates the scaling of Stark localization. The IPR for $\delta=-0.1$ across various system sizes is shown in Fig. 3(b1). Scaling analysis for IPR gives the value, $s=0.33$, which is confirmed by data collapse shown in Fig. 3(b2) and also by the cost-function approach. This scaling exponent is identical to that of the pure Stark model. In Fig. 3(c1), we plot the dependence of $\Delta E$ on $h$. The nature of the dependence is found to be the same as before. The collapse plot corresponding to $\Delta E$ is depicted in Fig. 3(c2), yielding a critical exponent of $z=2.00(2)$ which is again the same critical exponent value for that of the pure Stark model.
Thus, as we move away from the AA criticality the system does not experience much effect due to the AA potential. Likewise, the delocalization-localization transition is solely controlled by the Stark field. Therefore we study the behaviour of the scaling exponents as we slowly move away from the criticality. Fig. 4 nicely depicts this variation. The variation of $s/\nu$ and $z$ are shown in Fig. 4(a1) and Fig. 4(a2) respectively. To be specific, here, we only focus on the negative values of $\delta$, as we have found that a slight departure from the critical point in the positive $\delta$ values results in a very weird kind of variation in the characteristic observables and as a consequence of that, the scaling analysis fails there. This happens due to the competition between the two kinds of localization. From Fig. 4 it is easily noticeable that in the negative $\delta$ region, there is a continuous change in the values of the scaling exponents. As we move away from the AA criticality, the values of $s/\nu$ and $z$ slowly change from that of the pure AA model to the pure Stark model.
Moreover, in order to understand the competing nature of two distinct types of localization inducing potentials closely, it is interesting to directly look into the structures of the ground state wavefunctions for various cases. We present some exemplary cases in Fig. 5 for various scenarios relevant to our study of the scaling nature of the system near the criticality.
Figure 6: In (a1), we have plotted the curves of $\zeta L^{-1}$ against $hL^{1/\nu_{c}}$ for fixed system size, $L=377$, having various $\delta$ values: $\delta=-0.1$ (green circle), $\delta=-0.2$ (maroon square), $\delta=-0.3$ (blue diamond), $\delta=-0.4$ (orange triangle up), and $\delta=-0.5$ (black triangle left). Here $\nu_{c}=0.29$, the value of scaling exponent for localization length at the AA criticality, i.e., $\delta=0$. (a2) shows the collapse plot of $\zeta L^{-1}$ versus $hL^{1/\nu_{c}}(\delta L^{1/\nu_{\delta}})^{\kappa}$. Here also, we have taken $\phi$ average for 5000 choices.
VI Hybrid Scaling Exponent
Next, we focus on a case, where the system size, $L$, is kept fixed and different $\delta$ values are considered with $\delta<0$. For this case, we use the following scaling ansatz,
$\zeta/L=f_{7}(hL^{1/\nu_{c}}(\delta L^{1/\nu_{\delta}})^{\kappa}),$ (16)
where $\nu_{c}$ is the scaling exponent for the localization length of the ground state in our model, at the AA criticality, i.e., $\delta=0$. Here we numerically find out the new scaling exponent $\kappa$. Near AA criticality, this $\kappa$ should have the following form,
$\kappa=\nu_{\delta}(1/\nu_{s}-1/\nu_{c}),$ (17)
where $\nu_{\delta}$ and $\nu_{s}$ are the critical exponents for the localization length in the case of pure AA and Stark model.
In Fig. 6(a1), we show the variation of $\zeta/L$ against $hL^{1/\nu_{c}}$ for different values of $\delta$. Fig. 6(a2) shows that the collapse plot of $\zeta L^{-1}$ as a function of the hybrid quantity $hL^{1/\nu_{c}}(\delta L^{1/\nu_{\delta}})^{\kappa}$. From the cost function approach, we have found the scaling exponent $\kappa=-0.418(2)$ which matches with the value obtained via Eq. 17.
VII CONCLUSION
In this work we have explored the Stark localization in the presence of a quasi-periodic potential, namely the AA potential. This exploration is based on the scaling analysis of finite-size systems. First, we have studied the effect of increasing Stark field at the AA criticality. We have found that the scaling exponents $\nu$ and $s$ are quite different than both the pure AA model and the pure Stark model. However, the $s/\nu$ value is found to be the same as that of the pure AA model. Also, the scaling exponent $z$ is identical to that of the pure AA model. It indicates although the Stark field drives the transition, the transition is mostly AA type. Next, we have studied the same near the AA criticality. As in that case, we have used $\delta\neq 0$, we cannot employ the scaling ansatz for one control parameter-based phase transition. We have developed some scaling ansatz that is capable of extracting the relevant scaling exponents quite efficiently. Then, we take a fixed system size and vary the distance from the AA criticality. For this, a hybrid scaling ansatz is used and it successfully describes the scaling nature. Finally, we have looked into the variation of $s/\nu$ and $z$ with negative values of $\delta$ which have revealed the fact that near the AA criticality, these exponents coincide with that of the pure AA model, and as we move away from the AA criticality we enter into the region where the localization is solely driven by Stark field with almost no trace of the presence of AA potential.